# Beyond the Hodge Theorem: curl and asymmetric pseudodifferential projections

### v3

@@ Appendix E Taylor expansions for the operator of parallel transport
Recall that $Z_{\alpha}{}^{\beta}(x,y)$ is the map realising the parallel transport of vectors along the unique geodesic connecting $x$ to $y$, whereas $Z^{\alpha}{}_{\beta}(x,y)$ is the map realising the parallel transport of covectors along the same geodesic, see (4.6)–(4.8).
Proposition E.1.
@@
$Z_{\alpha}{}^{\beta}(0,y)=\delta_{\alpha}{}^{\beta}+\frac{1}{6}\operatorname{Riem}^{\beta}{}_{\mu\alpha\nu}(0)\,y^{\mu}y^{\nu}-\frac{1}{6}\frac{\partial^{2}\Gamma^{\beta}{}_{\mu\alpha}}{\partial y^{\nu}\partial y^{\rho}}(0)\,y^{\mu}y^{\nu}y^{\rho}+O(|y|^{4})\,$ (E.1)
@@
$Z^{\alpha}{}_{\beta}(0,y)=\delta^{\alpha}{}_{\beta}-\frac{1}{6}\operatorname{Riem}^{\alpha}{}_{\mu\beta\nu}(0)\,y^{\mu}y^{\nu}+\frac{1}{6}\frac{\partial^{2}\Gamma^{\alpha}{}_{\mu\beta}}{\partial y^{\nu}\partial y^{\rho}}(0)\,y^{\mu}y^{\nu}y^{\rho}+O(|y|^{4})\,.$ (E.2)
@@
$Z_{\alpha}{}^{\beta}(y,\tau y)=\delta_{\alpha}{}^{\beta}+\frac{\tau^{2}-1}{6}\operatorname{Riem}^{\beta}{}_{\mu\alpha\nu}(0)\,y^{\mu}y^{\nu}-\frac{\tau^{3}-1}{6}\frac{\partial^{2}\Gamma^{\beta}{}_{\mu\alpha}}{\partial y^{\nu}\partial y^{\rho}}(0)\,y^{\mu}y^{\nu}y^{\rho}+O(|y|^{4})\,$ (E.3)
@@
$Z^{\alpha}{}_{\beta}(y,\tau y)=\delta^{\alpha}{}_{\beta}-\frac{\tau^{2}-1}{6}\operatorname{Riem}^{\alpha}{}_{\mu\beta\nu}(0)\,y^{\mu}y^{\nu}+\frac{\tau^{3}-1}{6}\frac{\partial^{2}\Gamma^{\alpha}{}_{\mu\beta}}{\partial y^{\nu}\partial y^{\rho}}(0)\,y^{\mu}y^{\nu}y^{\rho}+O(|y|^{4})\,.$ (E.4)
Note that formula (E.1) agrees with [16, formula (7.1)].
Proof of Proposition E.1.
@@
$y^{\mu}\nabla_{\mu}v^{\beta}=y^{\mu}\left(\partial_{\mu}v^{\beta}+\Gamma^{\beta}{}_{\mu\alpha}v^{\alpha}\right)=0$ (E.5)
@@
$y^{\mu}\partial_{\mu}v^{\beta}=\dot{v}^{\beta}.$ (E.6)
Formulae (E.5) and (E.6) imply
[…]
@@
which follows from (E.5), (E.6) and the fact that Christoffel symbols vanish at the origin.
@@
$\frac{\partial\Gamma^{\beta}{}_{\mu\alpha}}{\partial y^{\nu}}(0)=-\frac{1}{3}\left(\operatorname{Riem}^{\beta}{}_{\mu\alpha\nu}(0)+\operatorname{Riem}^{\beta}{}_{\alpha\mu\nu}(0)\right)$
and performing integration in (E.7), we obtain
@@
The latter implies (E.1).
Combining (E.1) with (4.8) one immediately obtains (E.2).
In order to obtain (E.3) let us parallel transport along our straight line from $y$ to $\tau y$. Arguing as above, we get
@@
The latter implies (E.3) which, in turn, combined with (4.8) immediately gives us (E.4). ∎
Remark E.2.
For $\tau=0$ expansions (E.3) and (E.4) simplify to read
$Z_{\alpha}{}^{\beta}(y,0)=\delta_{\alpha}{}^{\beta}-\frac{1}{6}\operatorname{Riem}^{\beta}{}_{\mu\alpha\nu}(0)\,y^{\mu}y^{\nu}+\frac{1}{6}\frac{\partial^{2}\Gamma^{\beta}{}_{\mu\alpha}}{\partial y^{\nu}\partial y^{\rho}}(0)\,y^{\mu}y^{\nu}y^{\rho}+O(|y|^{4})\,$ (E.8)
@@
$Z^{\alpha}{}_{\beta}(y,0)=\delta^{\alpha}{}_{\beta}+\frac{1}{6}\operatorname{Riem}^{\alpha}{}_{\mu\beta\nu}(0)\,y^{\mu}y^{\nu}-\frac{1}{6}\frac{\partial^{2}\Gamma^{\alpha}{}_{\mu\beta}}{\partial y^{\nu}\partial y^{\rho}}(0)\,y^{\mu}y^{\nu}y^{\rho}+O(|y|^{4})\,.$ (E.9)
Formulae (E.1), (E.2), (E.8) and (E.9) agree with (4.7) and (4.8).
Appendix F An alternative derivation of formula (1.13)
@@
$g_{\alpha\beta}(x)=\delta_{\alpha\beta}-\frac{1}{6}(\nabla_{\sigma}\operatorname{Riem}_{\alpha\mu\beta\nu})(0)\,x^{\sigma}\,x^{\mu}\,x^{\nu}+O(|x|^{4}).$ (F.1)
@@
$[s]_{\alpha}{}^{\beta}(x,\xi)\sim\|\xi\|^{-1}\delta_{\alpha}{}^{\beta}+[s_{-2}]_{\alpha}{}^{\beta}(x,\xi)+[s_{-3}]_{\alpha}{}^{\beta}(x,\xi)+[s_{-4}]_{\alpha}{}^{\beta}(x,\xi)+\dots$
@@
$A_{\mathrm{prin}}(x,\xi)=-\varepsilon_{\beta}{}^{\alpha\gamma}\left(\,([s_{-3}]_{\alpha}{}^{\beta})_{x^{\gamma}}+i\,\xi_{\gamma}\,[s_{-4}]_{\alpha}{}^{\beta}\right)+O(|\xi|^{-3}|x|)\,.$ (F.2)
@@
In the remainder of this appendix, we will derive explicit formulae for $[s_{-3}]_{\alpha}{}^{\beta}$ and $[s_{-3}]_{\alpha}{}^{\beta}$. The proofs are straightforward, hence omitted.
Lemma F.2.
@@
$[q]_{\alpha}{}^{\beta}(x,\xi)\sim\|\xi\|^{2}\delta_{\alpha}{}^{\beta}+[q_{1}]_{\alpha}{}^{\beta}(x,\xi)+[q_{0}]_{\alpha}{}^{\beta}(x,\xi)$
be the (left) symbol of the operator $-\bm{\Delta}$. We have
$[q_{1}]_{\alpha}{}^{\beta}(x,\xi)=i\,a_{\alpha}{}^{\beta\gamma}{}_{\mu\nu}\,\xi_{\gamma}\,x^{\mu}\,x^{\nu}+O(|\xi|\,|x|^{3})\,,$ (F.3)
$[q_{0}]_{\alpha}{}^{\beta}(x)=b_{\alpha}{}^{\beta}{}_{\nu}\,x^{\nu}+O(|x|^{2})\,,$ (F.4)
@@
$a_{\alpha}{}^{\beta\gamma}{}_{\mu\nu}:=\left[\frac{1}{2}\nabla_{\mu}\operatorname{Ric}^{\gamma}{}_{\nu}-\frac{1}{12}\nabla^{\gamma}\operatorname{Ric}_{\mu\nu}\right]\delta_{\alpha}{}^{\beta}\\ -\frac{1}{6}\left[\nabla_{\alpha}\operatorname{Riem}^{\gamma}{}_{\mu}{}^{\beta}{}_{\nu}-3\nabla_{\mu}\operatorname{Riem}^{\gamma}{}_{\alpha}{}^{\beta}{}_{\nu}+5\nabla_{\nu}\operatorname{Riem}^{\gamma}{}_{\mu}{}^{\beta}{}_{\alpha}\right],$ (F.5)
$b_{\alpha}{}^{\beta}{}_{\nu}:=-\frac{1}{6}\nabla^{\beta}\operatorname{Ric}_{\alpha\nu}+\frac{1}{2}\nabla_{\alpha}\operatorname{Ric}^{\beta}{}_{\nu}+\frac{1}{2}\nabla_{\nu}\operatorname{Ric}_{\alpha}{}^{\beta}\,.$ (F.6)
Lemma F.3.
@@
$[r]_{\alpha}{}^{\beta}(x,\xi)\sim\|\xi\|\delta_{\alpha}{}^{\beta}+[r_{0}]_{\alpha}{}^{\beta}(x,\xi)+[r_{-1}]_{\alpha}{}^{\beta}(x,\xi)+[r_{-2}]_{\alpha}{}^{\beta}(x,\xi)+\dots$
and $s_{\alpha}{}^{\beta}$ of $(-\bm{\Delta})^{1/2}$ and $(-\bm{\Delta})^{-1/2}$ are expressed in terms of (F.3) and (F.4) via the following hierarchy of identities:
$[r_{0}]_{\alpha}{}^{\beta}=\frac{1}{2|\xi|}[q_{1}]_{\alpha}{}^{\beta}-\frac{1}{2i|\xi|^{2}}\xi^{\mu}(\|\xi\|)_{x^{\mu}}\,\delta_{\alpha}{}^{\beta}+O(|x|^{3})\,,$
$[r_{-1}]_{\alpha}{}^{\beta}\\ =\frac{1}{2|\xi|}[q_{0}]_{\alpha}{}^{\beta}-\frac{1}{2i|\xi|^{2}}\xi^{\mu}([r_{0}]_{\alpha}{}^{\beta})_{x^{\mu}}\\ +\frac{1}{4|\xi|}(|\xi|)_{\xi_{\mu}\xi_{\nu}}(\|\xi\|)_{x^{\mu}x^{\nu}}\,\delta_{\alpha}{}^{\beta}\\ +O(|\xi|^{-1}|x|^{2})\,,$
$[r_{-2}]_{\alpha}{}^{\beta}=-\frac{1}{2i|\xi|^{2}}\xi^{\mu}([r_{-1}]_{\alpha}{}^{\beta})_{x^{\mu}}\\ +\frac{1}{4|\xi|}(|\xi|)_{\xi_{\mu}\xi_{\nu}}([r_{0}]_{\alpha}{}^{\beta})_{x^{\alpha}x^{\beta}}+\frac{1}{12i|\xi|}(|\xi|)_{\xi_{\mu}\xi_{\nu}\xi_{\rho}}(\|\xi\|)_{x^{\mu}x^{\nu}x^{\rho}}\,\delta_{\alpha}{}^{\beta}\\ +O(|\xi|^{-2}|x|)\,,$
$[s_{-2}]_{\alpha}{}^{\beta}=-|\xi|^{-2}[r_{0}]_{\alpha}{}^{\beta}-\frac{1}{i|\xi|^{2}}\,\xi^{\mu}\,(\|\xi\|^{-1})_{x^{\mu}}\,\delta_{\alpha}{}^{\beta}+O(|\xi|^{-2}|x|^{3})\,,$
$[s_{-3}]_{\alpha}{}^{\beta}=-|\xi|^{-2}[r_{-1}]_{\alpha}{}^{\beta}-\frac{1}{i|\xi|^{2}}\xi^{\mu}\,([s_{-2}]_{\alpha}{}^{\beta})_{x^{\mu}}\\ +\frac{1}{2|\xi|}(|\xi|)_{\xi_{\mu}\xi_{\nu}}(\|\xi\|^{-1})_{x^{\mu}x^{\nu}}\,\delta_{\alpha}{}^{\beta}+O(|\xi|^{-3}|x|^{2})\,,$ (F.7)
$[s_{-4}]_{\alpha}{}^{\beta}=-|\xi|^{-2}[r_{-2}]_{\alpha}{}^{\beta}-\frac{1}{i|\xi|^{2}}\,\xi^{\mu}\,([s_{-3}]_{\alpha}{}^{\beta})_{x^{\mu}}\\ +\frac{1}{2|\xi|}(|\xi|)_{\xi_{\mu}\xi_{\nu}}([s_{-2}]_{\alpha}{}^{\beta})_{x^{\mu}x^{\nu}}+\frac{1}{6i|\xi|}(|\xi|)_{\xi_{\mu}\xi_{\nu}\xi_{\rho}}(\|\xi\|^{-1})_{x^{\mu}x^{\nu}x^{\rho}}\,\delta_{\alpha}{}^{\beta}\\ +O(|\xi|^{-4}|x|)\,.$ (F.8)
Remark F.4.
Lemmata F.2 and F.3 are true in any dimension $d$, not only in dimension $d=3$.
In order to verify our formula (1.13) for $A_{\mathrm{prin}}\,$, it only remains to compute (for example, with the help of computer algebra) the right-hand sides of (F.7) and (F.8), and substitute the resulting expressions into (F.2).
@@
$\operatorname{Ric}(x)=\begin{pmatrix}0&0&0\\ 0&0&x^{1}\\ 0&x^{1}&0\\ \end{pmatrix}+O(|x|^{2})$
(recall (6.24) and (F.1)) we get
$\left.s_{-3}\right|_{(x,\xi_{0})}=-\frac{1}{12}\begin{pmatrix}0&3x^{3}&3x^{2}\\ -x^{3}&0&3x^{1}\\ -x^{2}&3x^{1}&0\end{pmatrix}+O(|x|^{2}),\quad\left.s_{-4}\right|_{(x,\xi_{0})}=-\frac{i}{2}\begin{pmatrix}0&0&0\\ 1&0&0\\ 0&0&0\end{pmatrix}+O(|x|)\,,$ (F.9)
where $\xi_{0}$ is given by (6.26). Substituting (F.9) into (F.2) we obtain

### v4

@@ Appendix F Taylor expansions for the operator of parallel transport
[…]
Proposition F.1.
@@
$Z_{\alpha}{}^{\beta}(0,y)=\delta_{\alpha}{}^{\beta}+\frac{1}{6}\operatorname{Riem}^{\beta}{}_{\mu\alpha\nu}(0)\,y^{\mu}y^{\nu}-\frac{1}{6}\frac{\partial^{2}\Gamma^{\beta}{}_{\mu\alpha}}{\partial y^{\nu}\partial y^{\rho}}(0)\,y^{\mu}y^{\nu}y^{\rho}+O(|y|^{4})\,$ (F.1)
@@
$Z^{\alpha}{}_{\beta}(0,y)=\delta^{\alpha}{}_{\beta}-\frac{1}{6}\operatorname{Riem}^{\alpha}{}_{\mu\beta\nu}(0)\,y^{\mu}y^{\nu}+\frac{1}{6}\frac{\partial^{2}\Gamma^{\alpha}{}_{\mu\beta}}{\partial y^{\nu}\partial y^{\rho}}(0)\,y^{\mu}y^{\nu}y^{\rho}+O(|y|^{4})\,.$ (F.2)
@@
$Z_{\alpha}{}^{\beta}(y,\tau y)=\delta_{\alpha}{}^{\beta}+\frac{\tau^{2}-1}{6}\operatorname{Riem}^{\beta}{}_{\mu\alpha\nu}(0)\,y^{\mu}y^{\nu}-\frac{\tau^{3}-1}{6}\frac{\partial^{2}\Gamma^{\beta}{}_{\mu\alpha}}{\partial y^{\nu}\partial y^{\rho}}(0)\,y^{\mu}y^{\nu}y^{\rho}+O(|y|^{4})\,$ (F.3)
@@
$Z^{\alpha}{}_{\beta}(y,\tau y)=\delta^{\alpha}{}_{\beta}-\frac{\tau^{2}-1}{6}\operatorname{Riem}^{\alpha}{}_{\mu\beta\nu}(0)\,y^{\mu}y^{\nu}+\frac{\tau^{3}-1}{6}\frac{\partial^{2}\Gamma^{\alpha}{}_{\mu\beta}}{\partial y^{\nu}\partial y^{\rho}}(0)\,y^{\mu}y^{\nu}y^{\rho}+O(|y|^{4})\,.$ (F.4)
Note that formula (F.1) agrees with [17, formula (7.1)].
Proof of Proposition F.1.
@@
$y^{\mu}\nabla_{\mu}v^{\beta}=y^{\mu}\left(\partial_{\mu}v^{\beta}+\Gamma^{\beta}{}_{\mu\alpha}v^{\alpha}\right)=0$ (F.5)
@@
$y^{\mu}\partial_{\mu}v^{\beta}=\dot{v}^{\beta}.$ (F.6)
Formulae (F.5) and (F.6) imply
$\left.v^{\beta}\right|_{t=1}=\left.v^{\beta}\right|_{t=0}-\int_{0}^{1}y^{\mu}\Gamma^{\beta}{}_{\mu\gamma}(ty)\,v^{\gamma}(ty)\,dt\\ =\left.v^{\beta}\right|_{t=0}-\int_{0}^{1}y^{\mu}\left[\frac{\partial\Gamma^{\beta}{}_{\mu\alpha}}{\partial y^{\nu}}(0)\,y^{\nu}\,t+\frac{1}{2}\frac{\partial^{2}\Gamma^{\beta}{}_{\mu\alpha}}{\partial y^{\nu}\partial y^{\rho}}(0)\,y^{\nu}y^{\rho}\,t^{2}+O(|y|^{3})\right]\left[\left.v^{\alpha}\right|_{t=0}+O(|y|^{2})\right]\,dt\,.$ (F.7)
@@
which follows from (F.5), (F.6) and the fact that Christoffel symbols vanish at the origin.
@@
[…]
and performing integration in (F.7), we obtain
@@
The latter implies (F.1).
Combining (F.1) with (4.8) one immediately obtains (F.2).
In order to obtain (F.3) let us parallel transport along our straight line from $y$ to $\tau y$. Arguing as above, we get
@@
The latter implies (F.3) which, in turn, combined with (4.8) immediately gives us (F.4). ∎
Remark F.2.
For $\tau=0$ expansions (F.3) and (F.4) simplify to read
$Z_{\alpha}{}^{\beta}(y,0)=\delta_{\alpha}{}^{\beta}-\frac{1}{6}\operatorname{Riem}^{\beta}{}_{\mu\alpha\nu}(0)\,y^{\mu}y^{\nu}+\frac{1}{6}\frac{\partial^{2}\Gamma^{\beta}{}_{\mu\alpha}}{\partial y^{\nu}\partial y^{\rho}}(0)\,y^{\mu}y^{\nu}y^{\rho}+O(|y|^{4})\,$ (F.8)
@@
$Z^{\alpha}{}_{\beta}(y,0)=\delta^{\alpha}{}_{\beta}+\frac{1}{6}\operatorname{Riem}^{\alpha}{}_{\mu\beta\nu}(0)\,y^{\mu}y^{\nu}-\frac{1}{6}\frac{\partial^{2}\Gamma^{\alpha}{}_{\mu\beta}}{\partial y^{\nu}\partial y^{\rho}}(0)\,y^{\mu}y^{\nu}y^{\rho}+O(|y|^{4})\,.$ (F.9)
Formulae (F.1), (F.2), (F.8) and (F.9) agree with (4.7) and (4.8).
Appendix G An alternative derivation of formula (1.13)
@@
$g_{\alpha\beta}(x)=\delta_{\alpha\beta}-\frac{1}{6}(\nabla_{\sigma}\operatorname{Riem}_{\alpha\mu\beta\nu})(0)\,x^{\sigma}\,x^{\mu}\,x^{\nu}+O(|x|^{4}).$ (G.1)
@@
[…]
@@
$A_{\mathrm{prin}}(x,\xi)=-\varepsilon_{\beta}{}^{\alpha\gamma}\left(\,([s_{-3}]_{\alpha}{}^{\beta})_{x^{\gamma}}+i\,\xi_{\gamma}\,[s_{-4}]_{\alpha}{}^{\beta}\right)+O(|\xi|^{-3}|x|)\,.$ (G.2)
@@
In the remainder of this appendix, we will derive explicit formulae for $s_{-3}]_{\alpha}{}^{\beta}$ and $s_{-3}]_{\alpha}{}^{\beta}$. The proofs are straightforward, hence omitted.
Lemma G.2.
@@
[…]
be the (left) symbol of the operator $\bm{\Delta}$. We have
$q_{1}]_{\alpha}{}^{\beta}(x,\xi)=i\,a_{\alpha}{}^{\beta\gamma}{}_{\mu\nu}\,\xi_{\gamma}\,x^{\mu}\,x^{\nu}+O(|\xi|\,|x|^{3})\,$, (G.3)
$q_{0}]_{\alpha}{}^{\beta}(x)=b_{\alpha}{}^{\beta}{}_{\nu}\,x^{\nu}+O(|x|^{2})\,$, (G.4)
@@
$a_{\alpha}{}^{\beta\gamma}{}_{\mu\nu}:=\left[\frac{1}{2}\nabla_{\mu}\operatorname{Ric}^{\gamma}{}_{\nu}-\frac{1}{12}\nabla^{\gamma}\operatorname{Ric}_{\mu\nu}\right]\delta_{\alpha}{}^{\beta}\\ -\frac{1}{6}\left[\nabla_{\alpha}\operatorname{Riem}^{\gamma}{}_{\mu}{}^{\beta}{}_{\nu}-3\nabla_{\mu}\operatorname{Riem}^{\gamma}{}_{\alpha}{}^{\beta}{}_{\nu}+5\nabla_{\nu}\operatorname{Riem}^{\gamma}{}_{\mu}{}^{\beta}{}_{\alpha}\right$], (G.5)
$b_{\alpha}{}^{\beta}{}_{\nu}:=-\frac{1}{6}\nabla^{\beta}\operatorname{Ric}_{\alpha\nu}+\frac{1}{2}\nabla_{\alpha}\operatorname{Ric}^{\beta}{}_{\nu}+\frac{1}{2}\nabla_{\nu}\operatorname{Ric}_{\alpha}{}^{\beta}\,$. (G.6)
Lemma G.3.
@@
[…]
and $s_{\alpha}{}^{\beta}$ of $(-\bm{\Delta})^{1/2}$ and $(-\bm{\Delta})^{-1/2}$ are expressed in terms of (G.3) and (G.4) via the following hierarchy of identities:
$r_{0}]_{\alpha}{}^{\beta}=\frac{1}{2|\xi|}[q_{1}]_{\alpha}{}^{\beta}-\frac{1}{2i|\xi|^{2}}\xi^{\mu}(\|\xi\|)_{x^{\mu}}\,\delta_{\alpha}{}^{\beta}+O(|x|^{3})\,$,
$r_{-1}]_{\alpha}{}^{\beta}\\ =\frac{1}{2|\xi|}[q_{0}]_{\alpha}{}^{\beta}-\frac{1}{2i|\xi|^{2}}\xi^{\mu}([r_{0}]_{\alpha}{}^{\beta})_{x^{\mu}}\\ +\frac{1}{4|\xi|}(|\xi|)_{\xi_{\mu}\xi_{\nu}}(\|\xi\|)_{x^{\mu}x^{\nu}}\,\delta_{\alpha}{}^{\beta}\\ +O(|\xi|^{-1}|x|^{2})\,$,
$r_{-2}]_{\alpha}{}^{\beta}=-\frac{1}{2i|\xi|^{2}}\xi^{\mu}([r_{-1}]_{\alpha}{}^{\beta})_{x^{\mu}}\\ +\frac{1}{4|\xi|}(|\xi|)_{\xi_{\mu}\xi_{\nu}}([r_{0}]_{\alpha}{}^{\beta})_{x^{\alpha}x^{\beta}}+\frac{1}{12i|\xi|}(|\xi|)_{\xi_{\mu}\xi_{\nu}\xi_{\rho}}(\|\xi\|)_{x^{\mu}x^{\nu}x^{\rho}}\,\delta_{\alpha}{}^{\beta}\\ +O(|\xi|^{-2}|x|)\,$,
$s_{-2}]_{\alpha}{}^{\beta}=-|\xi|^{-2}[r_{0}]_{\alpha}{}^{\beta}-\frac{1}{i|\xi|^{2}}\,\xi^{\mu}\,(\|\xi\|^{-1})_{x^{\mu}}\,\delta_{\alpha}{}^{\beta}+O(|\xi|^{-2}|x|^{3})\,$,
$s_{-3}]_{\alpha}{}^{\beta}=-|\xi|^{-2}[r_{-1}]_{\alpha}{}^{\beta}-\frac{1}{i|\xi|^{2}}\xi^{\mu}\,([s_{-2}]_{\alpha}{}^{\beta})_{x^{\mu}}\\ +\frac{1}{2|\xi|}(|\xi|)_{\xi_{\mu}\xi_{\nu}}(\|\xi\|^{-1})_{x^{\mu}x^{\nu}}\,\delta_{\alpha}{}^{\beta}+O(|\xi|^{-3}|x|^{2})\,$, (G.7)
$s_{-4}]_{\alpha}{}^{\beta}=-|\xi|^{-2}[r_{-2}]_{\alpha}{}^{\beta}-\frac{1}{i|\xi|^{2}}\,\xi^{\mu}\,([s_{-3}]_{\alpha}{}^{\beta})_{x^{\mu}}\\ +\frac{1}{2|\xi|}(|\xi|)_{\xi_{\mu}\xi_{\nu}}([s_{-2}]_{\alpha}{}^{\beta})_{x^{\mu}x^{\nu}}+\frac{1}{6i|\xi|}(|\xi|)_{\xi_{\mu}\xi_{\nu}\xi_{\rho}}(\|\xi\|^{-1})_{x^{\mu}x^{\nu}x^{\rho}}\,\delta_{\alpha}{}^{\beta}\\ +O(|\xi|^{-4}|x|)\,.$ (G.8)
Remark G.4.
Lemmata G.2 and G.3 are true in any dimension $d$, not only in dimension $d=3$.
In order to verify our formula (1.13) for $A_{\mathrm{prin}}\,$, it only remains to compute (for example, with the help of computer algebra) the right-hand sides of (G.7) and (G.8), and substitute the resulting expressions into (G.2).
@@
$\operatorname{Ric}(x)=\begin{pmatrix}0&0&0\\ 0&0&x^{1}\\ 0&x^{1}&0\\ \end{pmatrix}+O(|x|^{2})\$
(recall (6.24) and (G.1)) we get
$\left.s_{-3}\right|_{(x,\xi_{0})}=-\frac{1}{12}\begin{pmatrix}0&3x^{3}&3x^{2}\\ -x^{3}&0&3x^{1}\\ -x^{2}&3x^{1}&0\end{pmatrix}+O(|x|^{2}),\quad\left.s_{-4}\right|_{(x,\xi_{0})}=-\frac{i}{2}\begin{pmatrix}0&0&0\\ 1&0&0\\ 0&0&0\end{pmatrix}+O(|x|)\,$, (G.9)
where $\xi_{0}$ is given by (6.26). Substituting (G.9) into (G.2) we obtain